jigglypuff314 (jigglypuff314):

Would the derivative of \[f(x) = x ^{3}(5x-2)^{2}\] be \[(x ^{2}(5x-2))(25x-6)\]?

OpenStudy (anonymous):

I would first exapand \(f(x)\) to make things a little easier: \[f(x)=x^3(5x-2)^2=x^3(25x^2-20x+4)=25x^5-20x^4+4x^3\] So then the derivative would be: \[f'(x)=125x^4-80x^3+12x^2\] Now you ended up with: \[f'(x)=x^2(5x-2)(25x-6)\] We can simplify: \[f'(x)=x^2(125x^2-30x-50x+12)=125x^4-80x^3+12x^2\] So in conclusion, YES! Haha

jigglypuff314 (jigglypuff314):

thank you! :)

OpenStudy (anonymous):

No prob Jigglypuff!